# Activation-Informed Merging of Large Language Models

### v2

@@ Abstract
Model merging, a method that combines the parameters and embeddings of multiple fine-tuned large language models (LLMs), offers a promising approach to enhance model performance across various tasks while maintaining computational efficiency. This paper introduces Activation-Informed Merging (AIM), a technique that integrates the information from the activation space of LLMs into the merging process to improve performance and robustness. AIM is designed as a flexible, complementary solution that is applicable to any existing merging method. It aims to preserve critical weights from the base model, drawing on principles from continual learning (CL) and model compression. Utilizing a task-agnostic calibration set, AIM selectively prioritizes essential weights during merging. We empirically demonstrate that AIM significantly enhances the performance of merged models across multiple benchmarks. Our findings suggest that considering the activation-space information can provide substantial advancements in the model merging strategies for LLMs with up to 40% increase in benchmark performance.
† Equal Contribution Authors
@@ 1 Introduction
Foundation models are rapidly becoming the dominant force for building Artificial Intelligence (AI) systems. In many cases, researchers build their machine learning models by starting from pre-trained foundation models and fine-tuning (FT) these pre-trained models for some desired target task (Pareja et al., 2024). In such a paradigm, numerous fine-tuned models are developed for various tasks. However, an important opportunity is missed, as these fine-tuned task-specialized models typically operate in isolation without leveraging the rich features that each possesses (Sudalairaj et al., 2024). This fact highlights the importance of a growing area of research focused on combining multiple task-specialized models fine-tuned from the same base foundation model.
@@
In particular, as large language models (LLMs) continue to evolve, it becomes increasingly important to develop methods that can effectively fuse the specialized knowledge of various fine-tuned models derived from the same foundation model. Model merging has shown broad applications, including enhancing accuracy and robustness (Wortsman et al., 2022), improving generalization (Ramé et al., 2023), multi-modal models (Sung et al., 2023), and model alignment to human feedback (Rame et al., 2024; Ramé et al., 2024). Given these benefits, a substantial amount of attention has been devoted to developing more effective merging algorithms for LLMs.
@@
In the vast majority of cases, merging LLMs is done using algorithms that explore the weight space of models and do not leverage the information in the activation space. Activation space information has been widely used to develop model pruning and compression methods, both in the context of general deep learning methods (Frantar and Alistarh, 2022), and more specifically for large language models Lin et al. (2024). However, this direction has remained under-explored for developing more robust merging algorithms. We hypothesize that the activation space information may hold key insights useful for model merging and explore this in our work.
@@
In this paper, we view the problem of merging from a continual learning perspective. Specifically, we explore how FT models and merging them can significantly deviate from the pre-trained base model, potentially leading to overall performance degradation. This is analogous to the common catastrophic forgetting problem in continual learning Farajtabar et al. (2020); Wang et al. (2024b); Sudalairaj et al. (2024). Given this perspective, we incorporate the activation space information, taking into account the importance of the base model in preserving its pre-trained capabilities while integrating new knowledge from fine-tuned models. To achieve this, we introduce a new method named Activation-Informed Merging (AIM), which modifies the update step in the merging process to ensure that the most influential weights of the base model, identified through its activations, undergo minimal changes. Figure 1 demonstrates this basic process of AIM.
@@
AIM fundamentally relates to the widely used approach of weight regularization in continual learning (Kirkpatrick et al., 2017; Ritter et al., 2018; Aljundi et al., 2018). When merging large language models fine-tuned from the same base model, the goal is to maintain the base model’s performance while improving the merged model’s expertise using the fine-tuned models. Various methods have been proposed to merge fine-tuned LLMs (Yu et al., 2024b, a; Ilharco et al., 2023; Wortsman et al., 2022; Davari and Belilovsky, 2024; Yadav et al., 2023; Matena and Raffel, 2022; Jin et al., 2023). Although useful, these methods are fragile to outlier and low-quality fine-tuned models and may perform worse than the base model. Hence, we take a new perspective toward merging and adopt the continual learning view to prevent catastrophic forgetting. Our extensive experimental study shows that AIM is a complementary solution that can be applied in conjunction with any prior merging methods and consistently improves their performances across all tested benchmarks—including math, code, and instruction following— by up to 40%. Despite its simplicity, AIM shows the effectiveness and importance of the activation space information for more effective model merging.
@@ 2 Background and Related Work
Fine-tuning pre-trained language models has become the prevalent paradigm for building downstream NLP models. Often, fine-tuned models are readily available, but their training data is not due to data privacy or intellectual property concerns (Dodge et al., 2020). With many fine-tuned checkpoints from the same pre-trained model, LLM merging has emerged as a complementary approach to fine-tuning, combining multiple task-specialized models into a unified architecture. This technique offers several advantages: it reduces storage and inference costs by consolidating expertise into a single model, facilitates compositional generalization by integrating specialized capabilities from different fine-tuned models, and supports decentralized model development by allowing independently trained models to be merged efficiently (Yadav et al., 2024).
@@
Model soup, proposed by Wortsman et al. (2022), demonstrates even simple averaging of the weights of multiple fine-tuned models can enhance accuracy and robustness in image classification and natural language processing applications. Further, Ramé et al. (2023) show that model merging can improve out-of-distribution generalization performance. Extending this approach beyond unimodal settings, Sung et al. (2023) empirically demonstrate that model merging is also effective in multimodal setups. Beyond generalization benefits, model merging has also been explored for alignment; WARP (Ramé et al., 2024) and WARM (Rame et al., 2024) introduce weight merging strategies to improve alignment in reinforcement learning from human feedback. WARP demonstrates that merging rewarded policies enhances model quality and alignment, while WARM shows that weight-averaged reward models improve robustness and help mitigate reward hacking.
@@
Recently, many methods have been proposed to go beyond simple weight averaging to merge fine-tuned LLMs into a multitask model by combining their capabilities. Spherical Linear intERPolation (SLERP) (Shoemake, 1985), originally proposed for animating rotations, which interpolates between two checkpoints, can be seen as a modification to simple weight averaging that finds a spherical path instead of a linear path in models’ parameter space. A main shortcoming of this approach is that it does not support merging more than two models. Jang et al. (2025) also leverages geometric insights, showing that merging only two fine-tuned models can provide superior in-distribution and out-of-distribution performance compared to ensembling multiple models.
@@
Task Arithmetic (Ilharco et al., 2023) generalizes simple weight averaging by introducing task vectors. It suggests moving the base model parameters in the direction of a weighted average of fine-tuned model differences with respect to the base model. Followed by the introduction of task vectors, Model Breadcrumbs (Davari and Belilovsky, 2024), Trim, Elect Sign & Merge (TIES merging) (Yadav et al., 2023), and Drop And REscale (DARE) (Yu et al., 2024a) leverage pruning techniques for better and more scalable ways of merging task vectors. WeIght DisENtanglement based merging (WIDEN) (Yu et al., 2024b) takes a more sophisticated approach to model merging by disentangling and analyzing weight components.
Another line of work on model merging takes advantage of the information in the model activations of the training data. Matena and Raffel (2022) propose Fisher merging, which leverages the Laplace approximation by using the diagonal of each model’s Fisher information. Jin et al. (2023) attempt to minimize prediction differences between the merged model and the individual models and introduce the Regression Mean (RegMean) method that calculates the optimal weights with respect to Euclidean distance to model predictions.
@@
We encourage the reader to refer to Yang et al. (2024) for a more comprehensive literature overview. The study provides a detailed discussion of model merging methods and theories, explores their applications in LLMs and multimodal large language models, and highlights future research directions.
@@
Continual learning strategies can be categorized into five overarching approaches: regularization, where parameters are constrained or updated based on past data; replay, where past data is replayed for the model as it encounters new data; optimization, where the loss function and optimizer are targeted; representation, where new data representations and learned embeddings can be exploited for less forgetting; and architecture, where models and parameters can expand as new data arrives Wang et al. (2024a). To avoid catastrophic forgetting of the base model’s abilities, we view the model merging problem through the lens of CL, primarily focusing on regularization-based methods. Regularization-based methods penalize deviation from the base model according to some norm (Shi et al., 2024). Various methods have been proposed to mitigate catastrophic forgetting such as Aljundi et al. (2018); Kirkpatrick et al. (2017); Ritter et al. (2018). In particular, Elastic Weight Consolidation (EWC) employs a Fisher Information Matrix to identify and protect parameters crucial for previous tasks by adding a quadratic penalty on the parameter shifts. Integrating CL approaches into the merging framework involves defining a weighted regularization term that selectively constrains parameter updates in critical areas for retaining previously learned tasks. This integration not only mitigates the risk of catastrophic forgetting but also enhances the adaptability and utility of the merged model.
@@
Using activation space information has been shown to be useful in the context of model compression. Frantar and Alistarh (2022) show that using a calibration dataset, deep learning models can be quantized and/or pruned efficiently. Lin et al. (2024) introduce Activation-aware Weight Quantization (AWQ) for LLM compression and show that protecting only 1% salient weights can greatly reduce quantization error.
@@ 3 Methodology: Activation-Informed Merging
As discussed in Section 2, most existing approaches for merging FT LLMs primarily focus on the weight space of the models being merged. However, it is well established that the activation space of these models contains crucial insights into the degree of importance of different parameters of LLMs. This was shown to be the case, for instance, in the work done by Lin et al. (2024) on Activation-aware Weight Quantization (AWQ), outperforming traditional quantization methods by including insight from the activation space of LLMs. Given this, we hypothesize that the activation space of LLMs likely holds useful clues for model merging as well. Inspired by AWQ, we introduce Activation-Informed Merging (AIM) for merging FT LLMs. In this section, we detail our proposed solution and discuss some of the inner workings of AIM.
@@ 3.1 The Merging Problem and Connections to Continual Learning
Consider the merging of models with parameters fine-tuned on different tasks from a common pre-trained model with parameters . For each fine-tuned LLM, we are essentially creating experts on specific tasks that move away from the generalist pre-trained model, hence usually degrading performance in some tasks while improving performance on the task for which the model is fine-tuned. In this sense, each FT model with parameters can be seen as a model fitted to a new task in a continual learning scenario with the potential for catastrophic forgetting on the generalist pre-trained model, which may not perform as well on the specific task but will have a more balanced performance across various tasks. As such, we hypothesize that when merging FT LLMs adapted from the same base model, emphasis on the base model can build better robustness to large performance degradation across numerous tasks while still allowing for capturing each FT expert’s capabilities. AIM seeks to achieve this by relaxing the changes to the salient weights of the base model in the final merged model. In this way, AIM is analogous to weight regularization in many continual learning approaches Wang et al. (2024b); Ritter et al. (2018); Kirkpatrick et al. (2017); Aljundi et al. (2018); Shi et al. (2024). Notably, the saliency of weights is determined by analyzing the activation space (sensitivities) of the base model rather than just regularizing based on the weight space, similar to Farajtabar et al. (2020); Min et al. (2022)
@@ 3.2 Activation Space Analysis
AIM determines the saliency of a model’s weights by looking at the scale of activations by passing a calibration dataset to the model and recording the scale of activations in each channel. To better understand why, we will analyze how perturbation of the weights of a given model affects the model outputs. Let the original weights be and the perturbation be , such that the perturbed weights are . The output of a linear layer with input and perturbed weights is:
@@
3.3 Calibration Data & Robustness to Calibration Data
We choose the calibration dataset to be a subset of the validation data from the pile dataset Gao et al. (2020), which is similar in distribution to most pre-training datasets. Most notably, this dataset has been utilized in model compression for both quantization in AWQ Lin et al. (2024) and pruning in WANDA Sun et al. (2024). This calibration dataset is considered to be fairly diverse and not task-specific, which should allow us to quantify weights’ saliency in a robust manner.
Size Doesn’t Matter
The use of activation spaces from calibration data has been thoroughly studied in model compression, by Lin et al. (2024) and Sun et al. (2024). In both cases, authors conduct extensive studies on the robustness of their methods, which, like ours, rely on the magnitude of activations, and both concur in two important findings. 1) When using calibration data to quantify activation space magnitudes to be used by algorithms for model quantization or pruning, performance is robust to dataset quality. This is in contrast to methods that require fine-tuning or retraining after compression. This observation is made clear in both studies and confirms that sensitivity to the quality of the data is much less in methods that require only activation space analysis without training. 2) Most notably, ablation studies on the size of the calibration set also show robustness to the size of calibration data, with both studies confirming that only 8-32 sample sequences of length approx. 2048 tokens are enough for model compression algorithms to produce robust outcomes Lin et al. (2024); Sun et al. (2024). Despite this robustness, we use the same 256 total sequences (approx. 524K tokens) that the authors of both studies use in their main experiments, however, this robustness to dataset size is noted as an important tool for reduction of computational cost and time for running algorithms such as ours. We also study this matter in an ablation study and show that like those studies AIM is also robust to dataset size and can be made much more efficient if need be.

### v3

@@ Abstract
Model merging, a method that combines the parameters and embeddings of multiple fine-tuned large language models (LLMs), offers a promising approach to enhance model performance across various tasks while maintaining computational efficiency. This paper introduces Activation-Informed Merging (AIM), a technique that integrates the information from the activation space of LLMs into the merging process to improve performance and robustness. AIM is designed as a flexible, complementary solution that is applicable to any existing merging method. It aims to preserve critical weights from the base model, drawing on principles from continual learning (CL) and model compression. Utilizing a task-agnostic calibration set, AIM selectively prioritizes essential weights during merging. We empirically demonstrate that AIM significantly enhances the performance of merged models across multiple benchmarks. Our findings suggest that considering the activation-space information can provide substantial advancements in the model merging strategies for LLMs with up to 40% increase in benchmark performance. Our code is publicly available at https://github.com/ahnobari/ActivationInformedMerging.
@@ 1 Introduction
Foundation models are rapidly becoming the dominant force for building Artificial Intelligence (AI) systems. In many cases, researchers build their machine learning models by starting from pre-trained foundation models and fine-tuning (FT) these pre-trained models for some desired target task [pareja2024]. In such a paradigm, numerous fine-tuned models are developed for various tasks. However, an important opportunity is missed, as these fine-tuned task-specialized models typically operate in isolation without leveraging the rich features that each possesses [sudalairaj2024lab]. This fact highlights the importance of a growing area of research focused on combining multiple task-specialized models fine-tuned from the same base foundation model.
@@
In particular, as large language models (LLMs) continue to evolve, it becomes increasingly important to develop methods that can effectively fuse the specialized knowledge of various fine-tuned models derived from the same foundation model. Model merging has shown broad applications, including enhancing accuracy and robustness [ModelSoups], improving generalization [rame2023ratatouille], multi-modal models [sung2023multimodalmerge], and model alignment to human feedback [WARM, WARP]. Given these benefits, a substantial amount of attention has been devoted to developing more effective merging algorithms for LLMs.
@@
In the vast majority of cases, merging LLMs is done using algorithms that explore the weight space of models and do not leverage the information in the activation space. Activation space information has been widely used to develop model pruning and compression methods, both in the context of general deep learning methods [frantar2022OBS], and more specifically for large language models [lin2024awqactivationawareweightquantization]. However, this direction has remained underexplored for developing more robust merging algorithms. We hypothesize that the activation space information may hold key insights useful for model merging and explore this in our work.
@@
In this paper, we view the problem of merging from a continual learning perspective. Specifically, we explore how FT models and merging them can significantly deviate from the pre-trained base model, potentially leading to overall performance degradation. This is analogous to the common catastrophic forgetting problem in continual learning [farajtabar2020OGD, Wang2024CLSurvey, sudalairaj2024lab]. Given this perspective, we incorporate the activation space information, taking into account the importance of the base model in preserving its pre-trained capabilities while integrating new knowledge from fine-tuned models. To achieve this, we introduce a new method named Activation-Informed Merging (AIM), which modifies the update step in the merging process to ensure that the most influential weights of the base model, identified through its activations, undergo minimal changes. Figure 1 demonstrates this basic process of AIM.
@@
AIM fundamentally relates to the widely used approach of weight regularization in continual learning [kirkpatrick2017overcomingCatastrophic, ritter2018laplaceCatastrophic, aljundi2018memory]. When merging large language models fine-tuned from the same base model, the goal is to maintain the base model’s performance while improving the merged model’s expertise using the fine-tuned models. Various methods have been proposed to merge fine-tuned LLMs [widen, DARE, TaskArithmetic, ModelSoups, davari2024modelbreadcrumbsscalingmultitask, yadav2023tiesmergingresolvinginterferencemerging, FisherMerging, jin2023dataless]. Although useful, these methods are fragile to outliers and low-quality fine-tuned models and may perform worse than the base model. Hence, we take a new perspective toward merging and adopt the continual learning view to prevent catastrophic forgetting. Our extensive experimental study shows that AIM is a complementary solution that can be applied in conjunction with any prior merging methods and consistently improves their performances across all tested benchmarks—including math, code, and instruction following—by up to 40%. Despite its simplicity, AIM shows the effectiveness and importance of the activation space information for more effective model merging.
@@ 2 Background and Related Work
Fine-tuning pre-trained language models has become the prevalent paradigm for building downstream NLP models. Often, fine-tuned models are readily available, but their training data is not due to data privacy or intellectual property concerns [dodge2020finetuningpretrainedlanguagemodels]. With many fine-tuned checkpoints from the same pre-trained model, LLM merging has emerged as a complementary approach to fine-tuning, combining multiple task-specialized models into a unified architecture. This technique offers several advantages: it reduces storage and inference costs by consolidating expertise into a single model, facilitates compositional generalization by integrating specialized capabilities from different fine-tuned models, and supports decentralized model development by allowing independently trained models to be merged efficiently [yadav2024mattersmodelmergingscale].
@@
Model soup, proposed by ModelSoups, demonstrates that even simple averaging of the weights of multiple fine-tuned models can enhance accuracy and robustness in image classification and natural language processing applications. Further, rame2023ratatouille show that model merging can improve out-of-distribution generalization performance. Extending this approach beyond unimodal settings, sung2023multimodalmerge empirically demonstrate that model merging is also effective in multimodal setups. Beyond generalization benefits, model merging has also been explored for alignment; WARP [WARP] and WARM [WARM] introduce weight merging strategies to improve alignment in reinforcement learning from human feedback. WARP demonstrates that merging rewarded policies enhances model quality and alignment, while WARM shows that weight-averaged reward models improve robustness and help mitigate reward hacking.
@@
Recently, many methods have been proposed to go beyond simple weight averaging to merge fine-tuned LLMs into a multitask model by combining their capabilities. Spherical Linear intERPolation (SLERP) [Slerp], originally proposed for animating rotations, which interpolates between two checkpoints, can be seen as a modification to simple weight averaging that finds a spherical path instead of a linear path in models’ parameter space. A main shortcoming of this approach is that it does not support merging more than two models. ModelStock also leverages geometric insights, showing that merging only two fine-tuned models can provide superior in-distribution and out-of-distribution performance compared to ensembling multiple models.
@@
Task Arithmetic [TaskArithmetic] generalizes simple weight averaging by introducing task vectors. It suggests moving the base model parameters in the direction of a weighted average of fine-tuned model differences with respect to the base model. Followed by the introduction of task vectors, Model Breadcrumbs [davari2024modelbreadcrumbsscalingmultitask], Trim, Elect Sign & Merge (TIES merging) [yadav2023tiesmergingresolvinginterferencemerging], and Drop And REscale (DARE) [DARE] leverage pruning techniques for better and more scalable ways of merging task vectors. WeIght DisENtanglement based merging (WIDEN) [widen] takes a more sophisticated approach to model merging by disentangling and analyzing weight components. Another line of work on model merging takes advantage of the information in the model activations of the training data. FisherMerging propose Fisher merging, which leverages the Laplace approximation by using the diagonal of each model’s Fisher information. jin2023dataless attempt to minimize prediction differences between the merged model and the individual models and introduce the Regression Mean (RegMean) method that calculates the optimal weights with respect to Euclidean distance to model predictions.
@@
We refer the reader to yang2024modelmergingllmsmllms for a more comprehensive literature overview. The study provides a detailed discussion of model merging methods and theories, explores their applications in LLMs and multimodal large language models, and highlights future research directions.
@@
Continual learning strategies can be categorized into five overarching approaches: regularization, where parameters are constrained or updated based on past data; replay, where past data is replayed for the model as it encounters new data; optimization, where the loss function and optimizer are targeted; representation, where new data representations and learned embeddings can be exploited for less forgetting; and architecture, where models and parameters can expand as new data arrives [Wang2024CLSurvey]. To avoid catastrophic forgetting of the base model’s abilities, we view the model merging problem through the lens of CL, primarily focusing on regularization-based methods. Regularization-based methods penalize deviation from the base model according to some norm [shi2024continualLLM]. Various methods have been proposed to mitigate catastrophic forgetting, e.g., aljundi2018memory, kirkpatrick2017overcomingCatastrophic, ritter2018laplaceCatastrophic. In particular, Elastic Weight Consolidation (EWC) employs a Fisher Information Matrix to identify and protect parameters crucial for previous tasks by adding a quadratic penalty on the parameter shifts. Integrating CL approaches into the merging framework involves defining a weighted regularization term that selectively constrains parameter updates in critical areas for retaining previously learned tasks. This integration not only mitigates the risk of catastrophic forgetting but also enhances the adaptability and utility of the merged model.
@@
Using activation space information has been shown to be useful in the context of model compression. frantar2022OBS show that using a calibration dataset, deep learning models can be quantized and/or pruned efficiently. [lin2024awqactivationawareweightquantization] introduce Activation-aware Weight Quantization (AWQ) for LLM compression and show that protecting only 1% salient weights can greatly reduce quantization error.
@@ 3 Methodology: Activation-Informed Merging
As discussed in Section 2, most existing approaches for merging FT LLMs primarily focus on the weight space of the models being merged. However, it is well established that the activation space of these models contains crucial insights into the degree of importance of different parameters of LLMs. This was shown to be the case, for instance, in the work done by lin2024awqactivationawareweightquantization on Activation-aware Weight Quantization (AWQ), outperforming traditional quantization methods by including insight from the activation space of LLMs. Given this, we hypothesize that the activation space of LLMs likely holds useful clues for model merging as well. Inspired by AWQ, we introduce Activation-Informed Merging (AIM) for merging FT LLMs. In this section, we detail our proposed solution and discuss some of the inner workings of AIM.
@@ 3.1 The Merging Problem and Connections to Continual Learning
Consider the merging of models with parameters fine-tuned on different tasks from a common pre-trained model with parameters . For each fine-tuned LLM, we are essentially creating experts on specific tasks that move away from the generalist pre-trained model, hence usually degrading performance in some tasks while improving performance on the task for which the model is fine-tuned. In this sense, each FT model with parameters can be seen as a model fitted to a new task in a continual learning scenario with the potential for catastrophic forgetting on the generalist pre-trained model, which may not perform as well on the specific task but will have a more balanced performance across various tasks. As such, we hypothesize that when merging FT LLMs adapted from the same base model, emphasis on the base model can build better robustness to large performance degradation across numerous tasks while still allowing for capturing each FT expert’s capabilities. AIM seeks to achieve this by relaxing the changes to the salient weights of the base model in the final merged model. In this way, AIM is analogous to weight regularization in many continual learning approaches [Wang2024CLSurvey, ritter2018laplaceCatastrophic, kirkpatrick2017overcomingCatastrophic, aljundi2018memory, shi2024continualLLM]. Notably, the saliency of weights is determined by analyzing the activation space (sensitivities) of the base model rather than just regularizing based on the weight space, similar to [farajtabar2020OGD, min2022one].
@@ 3.2 Activation Space Analysis
AIM determines the saliency of a model’s weights by looking at the scale of activations by passing a calibration dataset to the model and recording the scale of activations in each channel. To better understand why, we will analyze how perturbation of the weights of a given model affects the model outputs. Let the original weights be and the perturbation be , such that the perturbed weights are . The output of a linear layer with input and perturbed weights is
@@
3.3 Calibration Data and Robustness to Calibration Data
We choose the calibration dataset to be a subset of the validation data from the pile dataset [gao2020pile800gbdatasetdiverse], which is similar in distribution to most pre-training datasets. Most notably, this dataset has been utilized in model compression for both quantization in AWQ [lin2024awqactivationawareweightquantization] and pruning in WANDA [sun2024simpleeffectivepruningapproach]. This calibration dataset is considered to be fairly diverse and not task-specific, which should allow us to quantify weights’ saliency in a robust manner.
Robustness to Calibration Data
The use of activation spaces from calibration data has been thoroughly studied in model compression, by lin2024awqactivationawareweightquantization and sun2024simpleeffectivepruningapproach. In both cases, authors conduct extensive studies on the robustness of their methods, which, like ours, rely on the magnitude of activations, and both concur in two important findings. 1) When using calibration data to quantify activation space magnitudes to be used by algorithms for model quantization or pruning, performance is robust to dataset quality. This is in contrast to methods that require fine-tuning or retraining after compression. This observation is made clear in both studies and confirms that sensitivity to the quality of the data is much less in methods that require only activation space analysis without training. 2) Most notably, ablation studies on the size of the calibration set also show robustness to the size of calibration data, with both studies confirming that only 8-32 sample sequences of length approximately 2048 tokens are enough for model compression algorithms to produce robust outcomes [lin2024awqactivationawareweightquantization, sun2024simpleeffectivepruningapproach]. Despite this robustness, we use the same 256 total sequences (approximately 524K tokens) that the authors of both studies use in their main experiments; however, this robustness to dataset size is noted as an important tool for reducing computational cost and time for running algorithms such as ours. We also study this matter in an ablation study and show that, like those studies, AIM is also robust to dataset size and can be made much more efficient if need be.